Shift report 20000810 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Thu Aug 10 2000 - 07:31:14 EDT

  • Next message: Peter H. L. Christiansen: "Re: Shift report 20000810 00:00-08:00"

    ------------------------------------------------------------
    Date: Thu Aug 10 07:31:14 2000, Shift: 00:00-08:00
    Supervisor: HI, with: VS
    Log book#4 pages 44 - 44
    
    Shift Summary:
    ------------------------------------------------------------
    1.  1 short run (about 1 HR) run2181.  Since they tried with
    60 bounches(?), the intensity of beam was very high.
    Blue 6625, Yello 2781
    2.  There were tracks in all tpc.  2 Tiles for calibration 
    were read out.
    3.  H1 and C1 HV were off.  I do not know the reason.
    BRAVO does not easily communicates with those two HV.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Thu Aug 10 2000 - 07:34:46 EDT